OpenStudy (anonymous):

My problem is asking to solve the linear system by using substitution. the problem: 4x+3y=0 2x+y=-2

OpenStudy (anonymous):

Simple. Solve one equation for one of the variables. Substitute your result into the other equation and solve.

OpenStudy (anonymous):

rearrange second y= -2x-2 sub into first 4x +3 ( -2x-2) =0 -2x -6 =0 x=-3 sub back inot first y= 4 x=-3, y=4

OpenStudy (anonymous):

So, solving equation 2 for y gives y = -2x-2. Substitute for y in the first to get 4x + 3(-2x-2) = 0, so 4x - 6x -6 = 0, so -2x = 6, so x = -3.

OpenStudy (anonymous):

Then substitute -3 for x in either equation and solve for y.

OpenStudy (anonymous):

okayy. Got you. Thanx.